9"paper plate divided by 7.
How many inches around the perimeter to divide into 7 sections. Please help me with my daughter's 10th grade math. Thanks in advance- Ray

Found 2 solutions by mananth, edjones:
Answer by mananth(16949)

paper plate diemeter = 9"
pi* d = perimeter
3.14 * 9 = 28.26 inches
...
28.26/7 = 4.037 inches = 1 part
...
m.ananth@hotmail.ca

Answer by edjones(8007)

The diameter = 9"
Circumference=pi*d
=9*pi
=28.27...
28.27/7=4 inches around the perimeter (circumference) about.
.
Ed